A textbook publishing company has compiled data on total annual sales of its business texts for the preceding nine years:

Year 1 2 3 4 5 6 7 8 9
Sales (000): 37.20 44.80 47.40 51.60 55.70 57.30 60.50 69.90 73.80


a.
Using an appropriate model, forecast textbook sales for each of the next five years. (Round your intermediate calculations to 3 decimal places and final answers to 2 decimal places.)

Year 10 11 12 13 14
Forecast
Answer #1

a.

Using linear trend equation,

Year 10 11 12 13 14
Forecast 76.49 80.71 84.94 89.16 93.39

calculation

Year (x) Sales (000): (y) xy x2
1 37.2 37.2 1
2 44.8 89.6 4
3 47.4 142.2 9
4 51.6 206.4 16
5 55.7 278.5 25
6 57.3 343.8 36
7 60.5 423.5 49
8 69.9 559.2 64
9 73.8 664.2 81
Total 45 498.2 2744.6 285

x-bar = Sum(x)/n = 45/9 = 5
y-bar = Sum(y)/n = 498.2/9 = 55.35555556 = 55.356 (Rounded to 3 decimal places)

b = (Sum(xy) – n*x-bar*y-bar)/(Sum(x2) – n*x-bar*x-bar) = (2744.6-9*5*55.356)/(285–9*5*5)
   = 253.58/60= 4.226333333 = 4.226(Rounded to 3 decimal places)


a = y-bar –b*x-bar
= 55.356-4.226*5 = 34.226

Regression equation is y = a + bx,
we get following regression equation by substituting values of a and b
Regression equation is y =34.226 + 4.226x

Formula